OpenStudy (anonymous):

what is a quadratic function in the form of y = ax2 + bx + c that would have the given roots? 1. x = 2i and x = –2i 2. x =3 + i and x = 3 – i

OpenStudy (anonymous):

\[\LARGE x^2+px+q=0\] \[\LARGE x_1+x_2=-p \] \[\LARGE x_1\cdot x_2=q \]

OpenStudy (campbell_st):

expand and simplify (x - 2i)(x+2i) = 0 (x-(3+i))(x -(3-i))=0